OpenStudy (anonymous):

What is the solution set of |2x - 3| - 10 = -1?

OpenStudy (anonymous):

3y=-2x+6

OpenStudy (anonymous):

How did you get that @jonjenkins7653 ?

OpenStudy (anonymous):

y=-(2/3)x+2

OpenStudy (preetha):

Please provide an explanation and try to help the learner understand.

OpenStudy (anonymous):

We can't help him at all if we just give him answers and no explanation.

OpenStudy (anonymous):

I don't think that's right anyways.

OpenStudy (anonymous):

I learned that first you isolate the absolute value part of the equation and then set up two different equations so you can get your two possible x values. So first, add 10 to the both sides of the equation since that's the opposite of subtracting 10. Then your equation looks like this |2x - 3| = 9 Then you set up the two equations like this: 2x - 3 = 9 and 2x - 3 = -9 Then you solve both of them for x (which I assume you know how to do) and get your possible solutions. But I don't know if that's the answer you're looking for.

OpenStudy (anonymous):

x = 6 and x = -3

OpenStudy (anonymous):

Does that help?

OpenStudy (anonymous):

Yes it does, thankyou so much @ababylizard

OpenStudy (anonymous):

No problem. Glad to help. Good luck! :D